# Navigating Heterogeneity and Privacy in One-Shot Federated Learning with Diffusion Models

### v1

@@ Abstract
Federated learning (FL) enables multiple clients to train models collectively while preserving data privacy. However, FL faces challenges in terms of communication cost and data heterogeneity. One-shot federated learning has emerged as a solution by reducing communication rounds, improving efficiency, and providing better security against eavesdropping attacks. Nevertheless, data heterogeneity remains a significant challenge, impacting performance. This work explores the effectiveness of diffusion models in one-shot FL, demonstrating their applicability in addressing data heterogeneity and improving FL performance. Additionally, we investigate the utility of our diffusion model approach, FedDiff, compared to other one-shot FL methods under differential privacy (DP). Furthermore, to improve generated sample quality under DP settings, we propose a pragmatic Fourier Magnitude Filtering (FMF) method, enhancing the effectiveness of generated data for global model training.
@@ 1 Introduction
However, FL poses significant challenges in terms of communication cost and data heterogeneity across clients. Communication cost is a major bottleneck in FL systems, as clients need to communicate frequently with the server over multiple rounds during the training process [16, 28, 3]. This leads to a high communication overhead, making the process slow or simply infeasible. To overcome this challenge, one-shot federated learning has recently gained traction in the research community [10, 35, 36, 29]. In this setting, clients only communicate once with the server during the training process, significantly reducing the communication requirements. This approach not only improves the efficiency of the training process but also provides a better framework for privacy and application. Specifically, one-shot FL provides better security against eavesdropping attacks, where adversaries attempt to steal or tamper with the information being sent between clients and the server [22]. By only requiring one round of communication, one-shot FL significantly reduces the likelihood of such attacks. Furthermore, traditional multi-round training may not be a practical option in some cases, such as that of model markets [20]. In these scenarios, models are trained to convergence by a participating user, and simply made available as a pretrained model to potential buyers, without any option for iterative communication.
However, another significant challenge in federated learning still remains, and that is, the data heterogeneity problem [25, 17, 21, 16]. In FL, clients often have very different data distributions, making optimization particularly challenging across the federated system. In the one-shot setting, this is especially detrimental to performance. Without the luxury of multiple communication rounds, the resulting models will be significantly biased towards their narrow data distribution and difficult to reconcile into a global model. Knowledge distillation-based approaches have been studied in the literature in an attempt to address these problems [10, 20, 35]. Nonetheless, these methods still struggle immensely under high heterogeneity, resulting in large drops in performance.
@@
RQ1. First, we explore the utility of diffusion models in federated learning and their potential for improving the performance of the one-shot FL process. Diffusion models [14] have recently emerged as prominent approaches for image generation, inspiring our investigation. We suggest that specific traits of diffusion models could provide advantages for one-shot FL, as discussed in Section 3. We then validate this hypothesis through comprehensive experiments with our approach, FedDiff, across various settings.
RQ2. Second, we investigate one-shot FL methods under provable privacy budgets with differential privacy (DP), as this aspect is not addressed by existing state-of-the-art (SOTA) one-shot FL works. Safeguarding model privacy is critical in this setting, as the client models obtained in one-shot FL can be reused multiple times or even traded in a model market. Furthermore, in light of recent work [5], we examine the potential memorization of diffusion models within our FedDiff approach and the effectiveness of DP as a mitigation strategy.
After studying these research questions, we further explore a simple technique for improving the performance of our FedDiff method under DP settings. We observe that the quality of generated samples may deteriorate under DP constraints, rendering some samples counterproductive to the training of the global model. To improve the quality and consistency of the synthetic data, we propose a straightforward filtering approach, termed Fourier Magnitude Filtering (FMF). FMF leverages sample magnitudes derived from the Fourier transform to guide the selection of valuable samples. The resulting filtered dataset substantially improves the utility of the generated data, particularly in challenging conditions, as detailed in Section 5.3. Therefore, in this work, our contributions are summarized as follows:
  • We contribute to the FL literature with the first study exploring diffusion models in one-shot federated learning. Our comprehensive investigation unveils the unique advantages inherent to diffusion models, which enhances the overall performance of one-shot FL while also addressing the significant challenges of data heterogeneity. We therefore establish a novel approach, FedDiff, that not only ensures superior model performance but also aligns with the core requirements of one-shot FL.
  • We further study the privacy and utility of both discriminative and generative-based SOTA one-shot FL methods with DP guarantees under heterogeneous settings. We find that our FedDiff approach outperforms all other methods by a significant margin (from $\sim$5to $\sim$20across many datasets and settings), even when differential privacy is employed.
@@ 2 Background and Preliminaries
2.1 One-shot Federated Learning
Federated learning (FL) has emerged as a promising paradigm for collaborative machine learning across decentralized devices while preserving data privacy. The seminal work by McMahan et al. [24] introduced the concept of FL, where model updates are computed locally on user devices and aggregated on a central server. However, in the standard FL process, many iterative communication rounds are required for convergence. One-shot FL, therefore, studies how to effectively learn in this distributed setting in a single round, thereby mitigating the need for many communication rounds. Several approaches have been proposed to tackle the unique characteristics of one-shot FL. [10] introduce the one-shot federated learning framework and study several baseline approaches. In [10] and [20], distillation approaches are studied using the ensemble of client models to the global model, and assume a public dataset for this purpose. However, such an assumption is limited, as public data related to the domain of interest is often not available. A data-free method within the distillation methodology was proposed by [35], where a generative adversarial network (GAN) is trained at the server level to generate the data for distillation, and iteratively optimized between distilling to the server model and training the GAN with the ensemble of client models.
Nonetheless, these methods still struggle with heterogeneous environments, as we find in Section 3. Generative models on the client are well-suited for better undertaking in such settings, as they can focus on the narrow client distributions and simply generate data at the central location. [13] introduce the use of CVAEs in highly heterogeneous one-shot FL. However, CVAEs exhibit suboptimal sample quality, a limitation that becomes markedly exacerbated with more complex datasets and when subjected to the constraints of DP, which are not explicitly addressed in the study by [13]. In this work, we investigate diffusion models in one-shot FL and leverage their unique characteristics for the task, illustrating their potential in a variety of difficult FL settings and privacy guarantees.
@@ 2.1 One-shot Federated Learning
2.2 Diffusion Probabilistic Models
Diffusion probabilistic models [14, 6], or simply diffusion models as they are now commonly referenced (DM), have gained traction for application in generative vision tasks. Simply put, DMs aim to learn the backward process that can iteratively denoise an image corrupted with Gaussian noise back to the original. Specifically, as detailed in [14], noise is introduced to a given sample via a Markovian chain forward process
$q(x_{1:T}|x_{0})=\prod^{T}_{t=1}q(x_{t}|x_{t-1}),$ (1)
@@ 2.2 Diffusion Probabilistic Models
The process can also be conditioned on another variable $y$ in $\boldsymbol{\rho}_{\theta}\left(x_{t},y,t\right)$. For example, the diffusion model can be class conditioned [15], with $y$ being a variable representing the class of the sample from a classification dataset. We utilize the class-conditioning approach of [15] in our diffusion models for FL.
2.3 Differential Privacy
Differential privacy (DP) [8, 9, 7] is a framework for ensuring that the output of a computation, such as machine learning model training, does not reveal sensitive information about any individual data point in the training dataset. A computation is said to be differentially private if the probability of obtaining a particular output is roughly the same whether a particular individual’s data sample is included in the computation or not. Formally [9],
@@ 2.3 Differential Privacy
To train deep learning models with such guarantees, differentially private stochastic gradient descent (DPSGD) is typically employed [1]. In DPSGD, two main mechanisms are used to protect the privacy of individual data points: per-sample gradient clipping and the addition of random noise to the clipped gradients. Per-sample gradient clipping involves setting a maximum threshold on the norm of the gradient computed for each data point, so that if the norm of a gradient exceeds the threshold, it is rescaled. This step is necessary to limit the sensitivity of the loss function, which measures how much the loss function changes when a single data point is removed from the training dataset. After the gradients have been clipped, random noise is added to them before they are used to update the model parameters. The amount of noise added is calibrated based on privacy budget parameters $(\epsilon,\delta)$.
@@ 3 Diffusion Models for Federated Learning
Why diffusion models? In [32], a generative learning trilemma is shown with model types, trading off sample quality, diversity, and fast sampling. CVAEs (as employed in [13]) are typically identified to excel in diversity and fast sampling, but lacking in sample quality. However, for one-shot federated learning, fast sampling is not a concern, as the sampling can be done offline at the server (Figure 1). Therefore, high sample quality and diversity are more valuable properties in one-shot FL, as these will positively impact the performance of the trained global model with the synthetic data. In this trilemma, diffusion models excel in sample quality and diversity [32], but are not as quick to sample. This motivated us to investigate the potential of DMs in this setting, as the inherit strengths of DMs align with the needs of one-shot FL.
Furthermore, while CVAEs and diffusion models share a common origin in terms of their objective, they differ in their approach to achieving this objective. The optimization task of the diffusion model is simplified to learning a Markov process to reverse a fixed forward process. The training is structured such that the model only needs to learn how to denoise a small step in the generation process, breaking down the problem. In contrast, CVAEs must simultaneously learn both the forward process to encode the image to a latent space, and the decoding process from that latent vector. We reason that the simplified objective of DMs helps achieve superior performance when dealing with complex data within the challenging FL environment (data heterogeneity, class imbalance, and limited sample sizes). Moreover, in the FL setting, privacy is of critical importance. To ensure privacy, training is done with DP, which introduces noise to the training process and increases the difficulty of optimization. In these settings, the simpler training paradigm of diffusion models is potentially advantageous.
Overview. To provide a contextual foundation for our research inquiries, we start by laying out the settings of our study and approach in Section 3.1. With this groundwork, we investigate RQ1 in Section 4, where we delve into the effectiveness of diffusion models in one-shot FL with our FedDiff approach. In Section 5, we address RQ2 through a systematic exploration of one-shot FL methods within provable privacy budgets. Specifically, we evaluate FedDiff and other SOTA approaches under DP constraints, as well as investigate the viability of DP in mitigating memorization. In Section 5.3, we also introduce our Fourier Magnitude Filtering approach, aimed at enhancing the efficacy of generated data for global model training by selectively eliminating low-quality samples.
@@ 3.1 FedDiff and Experimental Setup
The basis of our approach, FedDiff, is illustrated in Figure 1. We begin by training class-conditioned diffusion models using the local data $\mathbf{x}$ on the clients. After training, the server collects these local models, denoted as $D_{0}$, $D_{1}$, …, $D_{c}$, which are then used to generate data $\mathbf{z}_{0}$, $\mathbf{z}_{1}$, …, $\mathbf{z}_{c}$. The label distributions from the clients are used to condition the generative models during generation, as in [13]. The combination of these synthesized samples forms our global training dataset, $\mathbf{G}$. Subsequently, the global model is trained on the synthetic dataset $\mathbf{G}$ and evaluated in our experiments.
@@
FedCVAE [13]. This recently proposed method employs conditional variational autoencoders (CVAEs) for one-shot federated learning. Their approach has two variants, FedCVAE-KD and FedCVAE-Ens, which differ in how they operate at the server level. FedCVAE-KD distills all generative models from the clients to a single CVAE, and then generates data for training the global model. On the other hand, FedCVAE-Ens employs each client model to generate data, contributing to the final dataset for training the server model. The latter variant always shows significantly better performance than the other in their paper; therefore, we compare with this FedCVAE-Ens variant and refer to it as FedCVAE in the rest of the paper.
Datasets. We employ three datasets, FashionMNIST [31], PathMNIST [33], and CIFAR-10 [18], which provide a range of domains and complexities. More details on the datasets are provided in the Supplementary Material. For our experiments, we divide the training set among $C$ clients with a Dirichlet distribution $Dir(\alpha)$, as commonly done in FL literature [25, 4, 12, 13]. This partitioning approach creates imbalanced subsets, where some clients may not have any samples for certain classes. As a result, a significant number of clients will only encounter a small subset (or potentially only one) of the available class instances. We visualize data distributions with $Dir(0.1)$ and $Dir(0.001)$ across 10 clients in the Supplementary Material.
Federated Learning Settings. We reproduce DENSE and FedCVAE for our settings with their respective official code repositories. For all experiments, we perform 3 independent runs with different seeds and report the mean and standard deviation. For all approaches, we train client models for 200 local epochs, as in [35]. For DENSE, FedCVAE, and FedDiff, we train the final global model for 50 epochs. For the generator of FedCVAE, we employ their CVAE variant with residual blocks, which has approximately 5.9M parameters. For our diffusion model, we employ a basic U-Net structure with residual blocks [14, 27] and class-conditioning, with similar parameters to FedCVAE ($\sim$5.8M). We employ a ResNet16 architecture for the discriminative models with approximately 6.4M parameters. For experiments with differential privacy, we employ the Opacus [34] library in PyTorch [26] to track privacy budgets. Further training details and additional experiments can be found in Supplementary Material.
4 RQ1: FedDiff for One-Shot FL
We investigate RQ1 by exploring the efficacy of our FedDiff approach and other SOTA one-shot FL methods across important FL scenarios, including different data heterogeneity levels, number of clients, and resource requirements.
Table 1: Data heterogeneity results with various $Dir(\alpha)$ partitions. Smaller alpha values indicate higher levels of heterogeneity. Typical approaches leveraging discriminative models rapidly degrade in performance as heterogeneity increases. However, generative approaches are more robust to such conditions. Our FedDiff shows superior performance to all, particularly in the most challenging scenarios (CIFAR-10, high heterogeneity).

### v2

@@ Abstract
Federated learning (FL) enables multiple clients to train models collectively while preserving data privacy. However, FL faces challenges in terms of communication cost and data heterogeneity. One-shot federated learning has emerged as a solution by reducing communication rounds, improving efficiency, and providing better security against eavesdropping attacks. Nevertheless, data heterogeneity remains a significant challenge, impacting performance. This work explores the effectiveness of diffusion models in one-shot FL, demonstrating their applicability in addressing data heterogeneity and improving FL performance. Additionally, we investigate the utility of our diffusion model approach, FedDiff, compared to other one-shot FL methods under differential privacy (DP). Furthermore, to improve generated sample quality under DP settings, we propose a pragmatic Fourier Magnitude Filtering (FMF) method, enhancing the effectiveness of the generated data for global model training. Code available at https://github.com/mmendiet/FedDiff.
@@ 1 Introduction
However, FL poses significant challenges in terms of communication cost and data heterogeneity across clients. Communication cost is a major bottleneck in FL systems, as clients need to communicate frequently with the server over multiple rounds during the training process [17, 30, 3]. This leads to a high communication overhead, making the process slow or infeasible. To overcome this challenge, one-shot FL has recently gained traction in the research community [11, 38, 39, 31]. In this setting, clients only communicate once with the server during the training process, significantly reducing the communication requirements. This approach not only improves the efficiency of the training process but also provides a better framework for privacy and application. Specifically, one-shot FL provides better security against eavesdropping attacks, where adversaries attempt to steal or tamper with the information being sent between clients and the server [23]. By only requiring one round of communication, one-shot FL significantly reduces the likelihood of such attacks. Furthermore, traditional multi-round training may not be a practical option in some cases, such as that of model markets [21]. In these scenarios, models are trained to be converged by a participating user and simply made available as a pretrained model to potential buyers without any option for iterative communication.
However, another significant challenge in federated learning still remains, and that is the data heterogeneity problem [26, 18, 22, 17]. In FL, clients often have very different data distributions, making optimization particularly challenging across the federated system. In the one-shot setting, this is especially detrimental to performance. Without the luxury of multiple communication rounds, the resulting models will be significantly biased towards their narrow data distribution and difficult to reconcile into a global model. Knowledge distillation-based approaches have been studied in the literature in an attempt to address these problems [11, 21, 38]. Nonetheless, these methods still struggle immensely under high heterogeneity, resulting in large drops in performance.
@@
RQ1. We explore the utility of diffusion models in FL and their potential for improving the performance of the one-shot FL process. Diffusion models (DMs) [15] have emerged as prominent approaches for image generation, inspiring our investigation. In the trilemma between sample quality, diversity, and inference speed of generative models [34], DMs excel in high sample quality and diversity. For one-shot FL, where inference occurs offline at the server, these two qualities where DMs excel are most crucial. Therefore, we suggest that these and other specific traits of DMs are well-suited and could provide advantages for one-shot FL, as further detailed in Sec. 3. We then validate this hypothesis through comprehensive experiments with our approach, FedDiff, across various settings.
RQ2. Generative models, while offering significant advantages for one-shot FL, can introduce privacy concerns. During generation, these models may produce images that are too similar to the original training data, potentially compromising data privacy. Therefore, we investigate one-shot FL methods under provable privacy budgets with differential privacy (DP), as this aspect is not addressed by existing state-of-the-art one-shot FL works. Safeguarding model privacy is critical in this setting since client models obtained in one-shot FL can be reused multiple times or traded in a model market. Furthermore, in light of recent work [5], we examine the potential memorization of diffusion models within our FedDiff approach and the effectiveness of DP as a mitigation strategy.
After studying these research questions, we further explore a simple technique for improving the performance of our FedDiff method under DP settings. We observe that the quality of generated samples may deteriorate under DP constraints, rendering some samples counterproductive to the training of the global model. To improve the quality and consistency of the synthetic data, we propose a straightforward filtering approach, termed Fourier Magnitude Filtering (FMF). FMF leverages sample magnitudes derived from the Fourier transform to guide the selection of valuable samples. The resulting filtered dataset substantially improves the utility of the generated data, particularly in challenging conditions, as detailed in Sec. 5.3. Therefore, in this work, our contributions are summarized as follows:
  • We contribute to the FL literature with the first study exploring diffusion models in one-shot federated learning. Our comprehensive investigation unveils the unique advantages inherent to diffusion models, enhancing the overall performance of one-shot FL while also addressing the significant challenges of data heterogeneity. We therefore establish a novel approach, FedDiff, that not only ensures superior model performance but also aligns with the core requirements of one-shot FL.
  • We study the privacy and utility of discriminative and generative-based SOTA one-shot FL methods with differential privacy (DP) guarantees under heterogeneous settings. Our FedDiff approach outperforms all other methods by a significant margin (from $\sim$5to $\sim$20across many datasets and settings), even when DP is employed.
@@ 2 Background and Preliminaries
One-shot Federated Learning. Federated learning (FL) has emerged as a promising paradigm for collaborative machine learning across decentralized devices while preserving data privacy. The seminal work by McMahan et al. [25] introduced the concept of FL, where model updates are computed locally on user devices and aggregated on a central server. However, in the standard FL process, many iterative communication rounds are required for convergence. One-shot FL studies how to effectively learn in this distributed setting in a single round, thereby mitigating the need for many communication rounds. Several approaches have been proposed to tackle the unique characteristics of one-shot FL. [11] introduce the one-shot federated learning framework and study several baseline approaches. In [11] and [21], distillation approaches are studied using the ensemble of client models to train the global model, assuming a public dataset for this purpose. However, such an assumption is limited, as public data related to the domain of interest is often not available. A data-free method within the distillation methodology was proposed by [38], where a generative adversarial network (GAN) is trained at the server level to generate the data for distillation, and iteratively optimized between distilling to the server model and training the GAN with the ensemble of client models.
Nonetheless, these methods still struggle with heterogeneous environments, as we find in Sec. 3. Generative models on the client are well-suited for better undertaking in such settings, as they can focus on the narrow client distributions and simply generate data at the central location. [14] introduce the use of CVAEs in highly heterogeneous one-shot FL. However, CVAEs exhibit suboptimal sample quality, a limitation that becomes markedly exacerbated with more complex datasets and when subjected to the constraints of DP, which are not explicitly addressed in the study by [14]. In this work, we investigate diffusion models in one-shot FL and leverage their unique characteristics for the task, illustrating their potential in a variety of difficult FL settings and privacy guarantees.
@@
Diffusion Probabilistic Models. Diffusion probabilistic models [15, 6], or simply diffusion models as they are now commonly referenced (DM), have gained traction for application in generative vision tasks. Simply put, DMs aim to learn the backward process that can iteratively denoise an image corrupted with Gaussian noise back to the original. Specifically, as detailed in [15], noise is introduced to a given sample via a Markovian chain forward process
$q(x_{1:T}|x_{0})=\prod^{T}_{t=1}q(x_{t}|x_{t-1}),\vspace{-2mm}$ (1)
@@
The process can also be conditioned on another variable $y$ in $\boldsymbol{\rho}_{\theta}\left(x_{t},y,t\right)$. For example, the diffusion model can be class conditioned [16], with $y$ being a variable representing the class of the sample from a classification dataset. We utilize the class-conditioning approach of [16] in our DMs for FL.
Differential Privacy. Differential privacy (DP) [8, 9, 7] is a framework for ensuring that the output of a computation, such as machine learning model training, does not reveal sensitive information about any individual data point in the training dataset. A computation is said to be differentially private if the probability of obtaining a particular output is roughly the same whether a particular individual’s data sample is included in the computation or not. Formally [9],
@@
To train deep learning models with such guarantees, differentially private stochastic gradient descent (DPSGD) is typically employed [1]. In DPSGD, two main mechanisms are used to protect the privacy of individual data points: per-sample gradient clipping and the addition of random noise to the clipped gradients. Per-sample gradient clipping involves setting a maximum threshold on the norm of the gradient computed for each data point so that, if the norm of a gradient exceeds the threshold, it is rescaled. This step is necessary to limit the sensitivity of the loss function, which measures how much the loss function changes when a single data point is removed from the training dataset. After the gradients have been clipped, random noise is added to them before they are used to update the model parameters. The amount of noise added is calibrated based on privacy budget parameters $(\epsilon,\delta)$.
@@ 3 Diffusion Models for Federated Learning
Why diffusion models? In [34], a generative learning trilemma is shown with model types, trading off sample quality, diversity, and fast sampling. CVAEs (as employed in [14]) are typically identified to excel in diversity and fast sampling, but lacking in sample quality. However, for one-shot federated learning, fast sampling is not a concern, as the sampling can be done offline at the server (Figure 1). Therefore, high sample quality and diversity are more valuable properties in one-shot FL, as these will positively impact the performance of the trained global model with the synthetic data. In this trilemma, diffusion models excel in sample quality and diversity [34], but are not as quick to sample. This motivated us to investigate the potential of DMs in this setting, as the inherent strengths of DMs align with the needs of one-shot FL.
Furthermore, while CVAEs and diffusion models share a common origin in terms of their objective, they differ in their approach to achieving this objective. The optimization task of the diffusion model is simplified to learning a Markov process to reverse a fixed forward process. The training is structured such that the model only needs to learn how to denoise a small step in the generation process, breaking down the problem. In contrast, CVAEs must simultaneously learn both the forward process to encode the image to a latent space and the decoding process from that latent vector. We reason that the simplified objective of DMs helps achieve superior performance when dealing with complex data within the challenging FL environment (data heterogeneity, class imbalance, and limited sample sizes). Moreover, in the FL setting, privacy is of critical importance. To ensure privacy, training is done with DP, which introduces noise to the training process and increases the difficulty of optimization. In these settings, the simpler training paradigm of diffusion models is potentially advantageous.
Overview. To provide a contextual foundation for our research inquiries, we start by laying out the settings of our study and approach in Sec. 3.1. With this groundwork, we investigate RQ1 in Sec. 4, where we delve into the effectiveness of diffusion models in one-shot FL with our FedDiff approach. In Sec. 5, we address RQ2 through a systematic exploration of one-shot FL methods within provable privacy budgets. We evaluate FedDiff and other SOTA approaches under DP constraints, as well as investigate the viability of DP in mitigating memorization. In Sec. 5.3, we introduce our Fourier Magnitude Filtering approach, aimed at enhancing the efficacy of generated data for global model training by selectively eliminating low-quality samples.
@@ 3.1 FedDiff and Experimental Setup
The basis of our FedDiff  is shown in Figure 1. We begin by training class-conditioned diffusion models using the local data $\mathbf{x}$ on the clients. After training, the server collects these local models, denoted as $D_{0}$, $D_{1}$, …, $D_{c}$, which are then used to generate data $\mathbf{z}_{0}$, $\mathbf{z}_{1}$, …, $\mathbf{z}_{c}$. The label distributions from the clients are used to condition the generative models during generation, as in [14]. The combination of these synthesized samples forms our global training dataset, $\mathbf{G}$. Subsequently, the global model is trained on the synthetic dataset $\mathbf{G}$ and evaluated in our experiments.
@@
FedCVAE [14]. This recently proposed method employs conditional variational autoencoders (CVAEs) for one-shot federated learning. Their approach has two variants, FedCVAE-KD and FedCVAE-Ens, which differ in how they operate at the server level. FedCVAE-KD distills all generative models from the clients to a single CVAE and then generates data for training the global model. On the other hand, FedCVAE-Ens employs each client model to generate data, contributing to the final dataset for training the server model. The latter variant always shows significantly better performance than the other in their paper; therefore, we compare with this FedCVAE-Ens variant and refer to it as FedCVAE in the rest of the paper.
Datasets. We employ three datasets, FashionMNIST [33], PathMNIST [35], and CIFAR-10 [19], which provide a range of domains and complexities. More details on the datasets are provided in the Supp. Material. For our experiments, we divide the training set among $C$ clients with a Dirichlet distribution $Dir(\alpha)$, as commonly done in FL literature [26, 4, 13, 14]. This partitioning approach creates imbalanced subsets, where some clients may not have any samples for certain classes. As a result, a significant number of clients will only encounter a small subset (or potentially only one) of the available class instances. We visualize data distributions with $Dir(0.1)$ and $Dir(0.001)$ across 10 clients in the Supp. Material.
Federated Learning Settings. We reproduce DENSE and FedCVAE for our settings with their respective official code repositories. For all experiments, we perform 3 independent runs with different seeds and report the mean and standard deviation. For all approaches, we train client models for 200 local epochs, as in [38]. For DENSE, FedCVAE, and FedDiff, we train the final global model for 50 epochs. For the generator of FedCVAE, we employ their CVAE variant with residual blocks, which has approximately 5.9M parameters. For our diffusion model, we employ a basic U-Net structure with residual blocks [15, 29] and class-conditioning, with similar parameters to FedCVAE ($\sim$5.8M). We employ a ResNet16 architecture for the discriminative models with approximately 6.4M parameters. For experiments with differential privacy, we employ the Opacus [37] library in PyTorch [28] to track privacy budgets. Further details can be found in Supp. Material.
Table 1: Data heterogeneity results with various $Dir(\alpha)$ partitions. Smaller $\alpha$ indicates higher levels of heterogeneity. Approaches leveraging discriminative models rapidly degrade in performance as heterogeneity increases. However, generative approaches are more robust to such conditions. Our FedDiff shows superior performance to all, particularly in the most challenging scenarios (CIFAR-10, high heterogeneity).
@@
(a) FedCVAE
(b) FedDiff (ours)
Figure 2: Random sets of generated samples from FedCVAE and our FedDiff. By leveraging the intrinsic properties of diffusion models, which are well-aligned with the requirements of one-shot FL, we achieve substantial benefits in sample quality and subsequent global model performance.
4 RQ1: FedDiff for One-Shot FL
We investigate RQ1 by exploring the efficacy of our FedDiff  and other SOTA one-shot FL methods across important FL scenarios, including different data heterogeneity levels, number of clients, and resource requirements.